OpenStudy (ksaimouli):

x-1/x^2-1 use long division

OpenStudy (anonymous):

\[\frac{x-1}{x^2-1}=\frac{x-1}{(x+1)(x-1)}=\frac{1}{x+1}\]

OpenStudy (ksaimouli):

satelite would u help me to draw as a graph plz